OpenStudy (anonymous):

can you help me solve this equation : -2x^2 + 5x - 12 = 0

OpenStudy (anonymous):

2x^2-5x+12=0 the answers are complex

OpenStudy (anonymous):

and can you just explain to me how to process please?

OpenStudy (anonymous):

\[-2x^2 + 5x - 12 = 0\] \[Δ=b^2-4ac=5^2-4*(-12)*(-2)=-71\] \[x_1,_2=(-b \pm i \sqrt{-Δ})/2a=(-5\pm i \sqrt{-(-71)})/2*(-2)= 5/4\pm \sqrt{71}/4 i\]